OpenStudy (anonymous):

What are the coordinates of the point at which the line passing through the points (1, -3) and (-2, 4) intersects the y-axis?

OpenStudy (popitree):

0,4

OpenStudy (anonymous):

How do u do this though?

OpenStudy (anonymous):

And the answer is (0,-2/3), I want to know how to get to that answer with showing work.

OpenStudy (popitree):

ya while calculating verbally i missed, let me explain.. equation of a straight line , y = mx+c; where m is slope and c is the intercept on y axis. now m = (y2-y1))/(x2-x1) so m = -7/3 y = -7/3x +c, now put any x,y combination on the equation and solve for c, it will come as -2/3